Please help me with this problem: (x-3)raised to the fraction power of 2/3 =25
Answer by ankor@dixie-net.com(15649)

Please help me with this problem: (x-3)raised to the fraction power of 2/3 =25
(x-3)^(2/3) = 25
We can write it as
(x-3)^(2/3) = 5^2
Find the square root of both sides and you have
(x-3)^(1/3) = 5
Cube both sides
x - 3 = 5^3
x - 3 = 125
x = 125 + 3
x = 128
;
:
:
You can check this on a calc: enter (128-3)^(2/3), the result: 25!